How do we show that W= {(x,-3x,2x)|x€R} is a subspace of R³?Also find a basis for subspace U of R³ which satisfies R³=W⊕U?

1
Expert's answer
2021-06-08T07:32:32-0400

Let us show that "W= \{(x,-3x,2x)|x\in\mathbb R\}" is a subspace of "\mathbb R^3". Let "\alpha, \beta\in\mathbb R,\ (x,-3x,2x), (y,-3y,2y)\in W." Then "\alpha (x,-3x,2x)+\beta (y,-3y,2y)= (\alpha x,-3\alpha x,2\alpha x)+(\beta y,-3\beta y,2\beta y)=
 (\alpha x+\beta y,-3\alpha x-3\beta y,2\alpha x+2\beta y)= (\alpha x+\beta y,-3(\alpha x+\beta y),2(\alpha x+\beta y))\in W."

Therefore, "W= \{(x,-3x,2x)|x\in\mathbb R\}" is a subspace of "\mathbb R^3". This subspace is of dimention 1 generated by "w=(1,-3,2)."


Consider the vectors "v=(0,2,3)" and "u=(-13,-3,2)." Taking into account that for the dot products we have "w\cdot v=0-6+6=0, w\cdot u=-13+9+4=0,\ v\cdot u=-6+6=0," we conclude that the subspace "U=\langle v,u\rangle" satisfies "\mathbb R^3=W⊕U", and "U" is of dimention 2 with basis consisting of vectors "v" and "u."